# ExAct: A Video-Language Benchmark for Expert Action Analysis

### v1

@@ 3.4 Stage IV: Final Expert Review and Validation
First, for each QA sample, several domain experts meticulously review the corresponding video clip and all five answer options. Their task is to select the option most consistent with the visual information presented in the video. After submitting their answer, the ground-truth label is revealed. The expert is then asked to verify three criteria: (1) whether the visual content of the video clip supports the correctness of the positive (i.e., correct) expert commentary; (2) whether any of the distractor commentaries also describe actions that are visible or valid within the video; and (3) whether all five candidate answer options are free from grammatical, logical, or instructional flaws. The sample will be removed if any of these criteria are not met. The second criterion is especially important because if a distractor candidate also describes something that appears in the video, the question becomes ill-posed due to multiple correct answers. For example, suppose that the positive expert commentary states, “The participant plants their non-kicking foot beside the ball,” while a distractor commentary states, “The participant plants their foot behind the ball.” In this case, both candidate commentaries are visually observable, which means that the expert cannot answer it using a single answer. We remove all such samples to avoid ambiguity.
This verification process is conducted by 16 experts, each assigned to one of the 11 activity categories based on their area of expertise. Each sample is reviewed by at least one qualified expert to ensure consistency and domain relevance. A screenshot of the annotation interface, along with additional implementation details, is provided in the supplementary material.
@@ 4 Experimental Setup
Baseline Models. To thoroughly assess the challenges posed by ExAct, we evaluate various state-of-the-art Video-Language Models (VLMs), including proprietary models: GPT-4o [20], Gemini 1.5 Pro [46], and open-source models: LLaVA-Video [55], LLaVA-OneVision [22], Qwen2.5-VL [3], VideoLLaMA [53], InternVL2.5 [9], and PerceptionLM [10]. These models vary significantly in architecture, training corpus, and modality integration strategies, offering a broad and representative basis to evaluate expert-level feedback capabilities.
@@ 5.1 Main Results
Gemini 1.5 Pro [20] 43.91 42.83 52.10 51.78 41.21 41.89 39.86
GPT-4o [20] 44.70 43.47 52.75 46.30 53.30 33.89 46.70
Open-source VLMs
PerceptionLM-8B [10] 24.65 24.22 28.16 25.75 22.53 22.95 26.42
VideoLLaMA3-7B [53] 26.38 26.64 23.30 29.32 26.65 23.79 27.79
InternVL2.5-78B [9] 33.48 31.93 36.57 33.70 37.91 32.00 34.62
LLaVA-OneVision-72B [22] 35.44 33.65 43.04 33.42 35.44 30.53 43.51
Qwen2.5-VL-72B-Instruct [3] 35.67 35.62 37.86 33.97 36.26 32.63 38.50
LLaVA-Video-72B [55] 41.58 41.81 42.72 44.11 32.42 38.74 48.52
Table 2: ExAct evaluation results (QA accuracy) across six diverse physical domains: Sports (Basketball, Soccer, Bouldering), Bike Repair, Cooking, Health (COVID-19 safety, CPR), Music (Guitar, Piano, Violin), and Dance. None of the methods achieves over 45% accuracy, indicating a significant room for improvement for future video-language models.
Figure 4: (a) Accuracy of the best-performing model (GPT-4o) compared to human experts and non-experts across domains. While GPT-4o achieves the highest accuracy among VLMs, it still falls short of human performance: experts consistently achieve over 80% accuracy in most domains, with soccer and bouldering reaching up to 95%. (b) Performance of VLMs on Good Execution (GE) and Tips for Improvement (TIPS) commentary categories. Models such as PerceptionLM, InternVL2.5, LLaVA-OneVision, and Qwen2.5-VL perform better on GE, while proprietary models like GPT-4o and Gemini 1.5 Pro show stronger performance on TIPS.
In Table 2, we report the performance of several state-of-the-art video-language models (VLMs) on our ExAct benchmark. Overall, all models perform poorly, with none exceeding 45% accuracy compared with 20% random-choice baseline. Among the evaluated models, GPT-4o achieves the highest overall accuracy at 44.70%, followed closely by Gemini 1.5 Pro at 43.91%. These two proprietary models outperform all others across all six domains. Notably, they achieve over 50% accuracy in the domains of Bike Repair, Cooking, and Health. These domains tend to rely more heavily on procedural knowledge, such as understanding sequences of actions and planning the next steps. In contrast, domains like Sports, Music, and Dance involve highly specialized, fine-grained physical movements that require detailed visual perception and precise temporal understanding.
We also observe that proprietary models (e.g., GPT-4o and Gemini) consistently outperform the best-performing open-source alternatives. This performance gap highlights the limitations of current public models in capturing the fine-grained, domain-specific reasoning required for expert-level skill understanding. Additionally, model size appears to be a significant factor in fine-grained video comprehension. Smaller models such as PerceptionLM-8B [10] and VideoLLaMA3-7B [53] achieve only 24.65% and 26.38% accuracy respectively—barely above random chance. In comparison, larger open-source models reach at least 33% accuracy, suggesting that scale contributes to more effective representation and complex activity analysis.
@@ 5.2 Human Performance Analysis
To quantify the gap between human and model performance, we conduct a human evaluation involving two groups: experts, comprising trained coaches and professionals with significant domain expertise (i.e., $>10\ years$), and non-experts, individuals without professional expertise in a given domain. As shown in Figure 4 (a), experts achieve consistently high accuracy—often exceeding 80% across most domains, with soccer and bouldering reaching around 95%. Additionally, we observe that while non-experts perform considerably worse (60-70% accuracy), they still significantly surpass the best-performing VLM model (i.e., GPT-4o). These results highlight a substantial gap between human and VLM capabilities, emphasizing the challenges of ExAct and the limitations of current VLMs.
@@ 5.3 Performance by Expert Commentary Type
Figure 4 (b) shows the performance of various VLMs across two expert commentary categories: Good Execution (GE) and Tips for Improvement (TIPS). Models such as PerceptionLM [10], InternVL2.5 [9], LLaVA-OneVision [22], and Qwen2.5-VL [3] perform better on GE samples. However, they struggle with questions from the TIPS category, which require identifying subtle mistakes in skilled activity executions. In contrast, stronger models such as GPT-4o [20] and Gemini 1.5 Pro [46] perform better on questions from the TIPS category, indicating a greater capacity for expert-level understanding of errors/mistakes in physical human activities.
@@ 5.4 Ablation Studies
In Table 3, we conduct ablation studies on three key design choices: (a) number of input frames, (b) impact of LLM size, (c) different prompt strategies. For (a)–(c), we use LLaVA-Video [55], the strongest-performing open-source baseline. Additionally, we ablate the effect of spatial video resolution using Qwen2.5-VL, which supports native-resolution inputs.
@@
w/o Time Instruct. 40.04
w/ Time Instruct. 41.58
(c) Prompting strategies: Incorporating time instruction (Time Instruct.) into the prompt improves accuracy.
@@
Prompting Strategies. In Table 3, we analyze the impact of different prompting strategies on model performance. We first evaluate the role of including time instructions (e.g.“The video is {video_time}s long, and {len(frames)} uniformly sampled frames occur at {frame_time}.” in the prompt. Removing the time instruction leads to a performance degradation of 1.54%, indicating that the understanding of time plays a meaningful role in the model’s analysis of human skill.
Spatial Video Resolution. We investigate how different spatial input resolutions affect the performance of Qwen2.5-VL by evaluating three settings: the original resolution of 796$\times$448 (35.67% accuracy), a 1.5$\times$ downsampled version at 531$\times$299 (37.40% accuracy), and a 2$\times$ downsampled version at 398$\times$224 (35.03% accuracy). Interestingly, the model achieves the highest accuracy at the mid-level resolution of 531$\times$299. We hypothesize that although higher resolutions provide more visual detail, they may lead to performance degradation due to increased token length and a mismatch with the lower-resolution inputs commonly seen during pretraining.

### v2

@@ 3.4 Stage IV: Final Expert Review and Validation
First, for each QA sample, domain experts meticulously review the corresponding video clip and all five answer options. Their task is to select the option most consistent with the visual information presented in the video. After submitting their answer, the ground-truth label is revealed. The expert is then asked to verify three criteria: (1) whether the visual content of the video clip supports the correctness of the positive (i.e., correct) expert commentary; (2) whether any of the distractor commentaries also describe actions that are visible or valid within the video; and (3) whether all five candidate answer options are free from grammatical, logical, or instructional flaws. The sample will be removed if any of these criteria are not met. The second criterion is especially important because if a distractor candidate also describes something that appears in the video, the question becomes ill-posed due to multiple correct answers. For example, suppose that the positive expert commentary states, “The participant plants their non-kicking foot beside the ball,” while a distractor commentary states, “The participant plants their foot behind the ball.” In this case, both candidate commentaries are visually observable, which means that the expert cannot answer it using a single answer. We remove all such samples to avoid ambiguity.
This verification process is conducted by 16 experts, each assigned to one of the 11 activity categories based on their area of expertise. Each sample is reviewed by at least one qualified expert to ensure consistency and domain relevance. For domains reviewed by two experts, we observed that experts reached agreement in over 90% of cases—demonstrating strong consistency in their judgments. A screenshot of the annotation interface, along with additional implementation details, is provided in the supplementary material.
@@ 4 Experimental Setup
Baseline Models. To thoroughly assess the challenges posed by ExAct, we evaluate various state-of-the-art Video-Language Models (VLMs), including proprietary models: Gemini 2.5 Pro [comanici2025gemini], GPT-4.1 [hurst2024gpt], GPT-4o [hurst2024gpt], Gemini 1.5 Pro [team2024gemini], and open-source models: LLaVA-Video [zhang2024video], LLaVA-OneVision [li2024llava], Qwen2.5-VL [bai2025qwen2], VideoLLaMA [zhang2025videollama], InternVL2.5 [chen2024internvl], and PerceptionLM [cho2025perceptionlm]. These models vary significantly in architecture, training corpus, and modality integration strategies, offering a broad and representative basis to evaluate expert-level feedback capabilities.
@@ 5.1 Main Results
Open-source VLMs
PerceptionLM-8B [cho2025perceptionlm] 24.65 24.22 28.16 25.75 22.53 22.95 26.42
VideoLLaMA3-7B [zhang2025videollama] 26.38 26.64 23.30 29.32 26.65 23.79 27.79
InternVL2.5-78B [chen2024internvl] 33.48 31.93 36.57 33.70 37.91 32.00 34.62
LLaVA-OneVision-72B [li2024llava] 35.44 33.65 43.04 33.42 35.44 30.53 43.51
Qwen2.5-VL-72B-Instruct [bai2025qwen2] 35.67 35.62 37.86 33.97 36.26 32.63 38.50
LLaVA-Video-72B [zhang2024video] 41.58 41.81 42.72 44.11 32.42 38.74 48.52
@@
Gemini 1.5 Pro [team2024gemini] 43.91 42.83 52.10 51.78 41.21 41.89 39.86
GPT-4o [hurst2024gpt] 44.70 43.47 52.75 46.30 53.30 33.89 46.70
GPT-4.1 [hurst2024gpt] 50.89 51.37 58.90 54.25 51.10 40.84 51.48
Gemini 2.5 Pro [comanici2025gemini] 55.35 52.58 65.05 58.36 60.71 53.05 53.98
Table 2: ExAct evaluation results (QA accuracy) across six diverse physical domains: Sports (Basketball, Soccer, Bouldering), Bike Repair, Cooking, Health (COVID-19 safety, CPR), Music (Guitar, Piano, Violin), and Dance. The results show a significant gap between the performance of modern Vision-Language Models (VLMs) and human experts, indicating that there is a significant room for improvement for future video-language models.
Figure 4: (a) Accuracy of the best-performing model (Gemini 2.5 Pro) compared to human experts and non-experts across domains. While Gemini 2.5 Pro achieves the highest accuracy among VLMs, it still falls short of human performance: experts consistently achieve over 80% accuracy in most domains, with soccer and bouldering reaching up to 95%. (b) Performance of VLMs on Good Execution (GE) and Tips for Improvement (TIPS) commentary categories. Models such as PerceptionLM, InternVL2.5, LLaVA-OneVision, and Qwen2.5-VL perform better on GE, while all proprietary models show stronger performance on TIPS.
In Table 2, we report the performance of several state-of-the-art video-language models (VLMs) on our ExAct benchmark. Overall, none of the methods achieve accuracy exceeding 56%. Among the evaluated models, Gemini 2.5 Pro achieves the highest overall accuracy at 55.35%. This model outperforms all others across all six domains. Notably, it achieves over 60% accuracy in the domains of Bike Repair and Health. These domains tend to rely more heavily on procedural knowledge, such as understanding sequences of actions and planning the next steps. In contrast, domains like Sports, Cooking, Music, and Dance involve highly specialized, fine-grained physical movements that require detailed visual perception and precise temporal understanding.
We also observe that proprietary models (e.g., GPT and Gemini) consistently outperform the best-performing open-source alternatives. This performance gap highlights the limitations of current public models in capturing the fine-grained, domain-specific reasoning required for expert-level skill understanding. Additionally, model size appears to be a significant factor in fine-grained video comprehension. Smaller models such as PerceptionLM-8B [cho2025perceptionlm] and VideoLLaMA3-7B [zhang2025videollama] achieve only 24.65% and 26.38% accuracy respectively—barely above random chance. In comparison, larger open-source models reach at least 33% accuracy, suggesting that scale contributes to more effective representation and complex activity analysis.
@@ 5.2 Human Performance Analysis
To quantify the gap between human and model performance, we conduct a human evaluation involving two groups: experts, comprising trained coaches and professionals with significant domain expertise (i.e., >10years>10\ years), and non-experts, individuals without professional expertise in a given domain. As shown in Figure 4 (a), experts achieve consistently high accuracy—often exceeding 80% across most domains, with soccer and bouldering reaching around 95%. Additionally, we observe that while non-experts perform considerably worse (60-70% accuracy), they still surpass the best-performing VLM model (i.e., Gemini 2.5 Pro). These results highlight a substantial gap between human and VLM capabilities, emphasizing the challenges of ExAct and the limitations of current VLMs.
@@ 5.3 Performance by Expert Commentary Type
Figure 4 (b) shows the performance of various VLMs across two expert commentary categories: Good Execution (GE) and Tips for Improvement (TIPS). Models such as PerceptionLM [cho2025perceptionlm], InternVL2.5 [chen2024internvl], LLaVA-OneVision [li2024llava], and Qwen2.5-VL [bai2025qwen2] perform better on GE samples. However, they struggle with questions from the TIPS category, which require identifying subtle mistakes in skilled activity executions. In contrast, stronger models such as Gemini 2.5 Pro [comanici2025gemini] and GPT-4.1 [hurst2024gpt] perform better on questions from the TIPS category, indicating a greater capacity for expert-level understanding of errors/mistakes in physical human activities.
@@ 5.4 Ablation Studies
In Table 3, we conduct ablation studies on four key design choices: (a) number of input frames, (b) impact of LLM size, (c) different prompt strategies, and (d) spatial video resolution. For (a)–(c), we use LLaVA-Video [zhang2024video], the strongest-performing open-source baseline. We ablate the effect of spatial video resolution using Qwen2.5-VL, which supports native-resolution inputs.
@@
w/o Time Info 40.04
w/ Time Info 41.58
(c) Prompting strategies: Incorporating time information into the prompt improves accuracy.
@@
Prompting Strategies. In Table 3, we analyze the impact of different prompting strategies on model performance. We first evaluate the role of including time information (e.g.“The video is {video_time}s long, and {len(frames)} uniformly sampled frames occur at {frame_time}.” in the prompt. Removing the time information leads to a performance degradation of 1.54%, indicating that the understanding of time plays a meaningful role in the model’s analysis of human skill.
[…]
@@ 6 Conclusion
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the key contributions of the paper, including “ExAct will be beneficial for developing and evaluating VLMs capable of precise understanding of human skills in various physical and procedural domains.” These claims are consistent with the technical content and experimental results presented in the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We provide limitations in Section 6.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include any theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: All necessary details to reproduce the main experimental results are provided in Section 4.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We provide open access to both the dataset and code used in our experiments. The dataset is available at: https://huggingface.co/datasets/Alexhimself/ExAct, and the evaluation code is available at: https://github.com/Texaser/Exact. Detailed instructions for reproducing the main experimental results are included in the repository and supplemental material.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We provide the details in Section 4.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars. All results come from a single deterministic inference pass with pre-trained models—no model training or other stochastic procedures are involved—so repeated runs would yield identical numbers, while retraining the large models to obtain multiple seeds would be prohibitively expensive.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We provide the details in Section 4.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: Our research adheres fully to the NeurIPS Code of Ethics. We ensured responsible use of data, preserved participant anonymity where applicable, and followed ethical practices in the collection, use, and reporting of all results. No known risks, harms, or ethical violations were introduced by this work.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: The paper discusses both potential positive and negative societal impacts. It highlights the benefits of the proposed benchmark for evaluating expert-level understanding and supporting applications in education and skill training. In the limitations section, we also acknowledge potential negative societal impacts. Although all participants consented to data usage, the inclusion of real-world videos containing identifiable human faces may still pose privacy concerns.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The benchmark does not involve the release of models or datasets with a high risk of misuse. All data sources are curated from publicly available, non-sensitive content with appropriate usage rights, and the benchmark is intended for academic research under a permissive license.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: All external assets used in this work—including datasets, models, and code—are properly cited in the paper. Their respective licenses and terms of use (e.g., MIT, CC-BY, or custom academic licenses) have been reviewed and respected. We ensure that all assets are used in accordance with their intended academic and research use permissions.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: The paper introduces a new benchmark, and we provide detailed documentation alongside the released assets. This includes task definitions, annotation guidelines, dataset statistics, data format specifications, and instructions for evaluation and usage. The documentation is provided alongside the released code and dataset at the uploaded URL.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [Yes]
  69. Justification: As the paper involves research with human annotators, we provide the full text of the instructions given to participants, example screenshots of the annotation interface in the supplemental material.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [Yes]
  74. Justification: The study involves human annotators evaluating non-anonymized video clips. While the videos may contain identifiable individuals, the annotation task itself poses no known risks to participants. All annotators were informed of the task content and compensation before participation. Based on our institutional guidelines, the nature of the task did not require formal IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [Yes]
  79. Justification: Large language models (LLMs) are a core component of our evaluation methodology. Their usage is clearly documented in the main paper, including the specific models employed (e.g., GPT-4o, Gemini 1.5 Pro) and their roles within the evaluation pipeline.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
ExAct: A Video-Language Benchmark
for Expert Action Analysis